Use a double angle formula to rewrite each trigonometric ratio

a) sin (4a)
b) cos (6a)
c) tan (5a)
1
Expert's answer
2020-10-25T18:41:16-0400

(a) Sin(4a)="Sin2(2a)=2sin2acos2a"

Further It can be resolved as

="2\times 2sinacosa\times (2cos^2a-1)"


="8sinacos^3a-4sinacosa"


(b) Cos(6A)=cos2(3a)=2"cos^23a-1" or "1-2sin^23a"


(c) tan(5a)=tan2(2.5a)="\dfrac{2tan(2.5a)}{1-tan^22.5a}"